# Learning Selection Cuts With Gradients

### v1

@@ 6 Results
We first study the behavior of several network configurations using the Gaussian dataset. We consider: a traditional linear network with a BCE loss function; a one-to-one linear network with a BCE loss function; a one-to-one linear network with an efficiency loss function; and an ensemble of one-to-one linear networks trained at different efficiency working points. We then study the ability of these different networks to classify simulated LHC events as arising from Standard Model interactions or from production of new particles in the Minimal Supersymmetric Standard Model (MSSM).
6.1 Gaussian dataset
The simple linear network described in section 4 is trained using the Gaussian dataset, with results shown in Figure 7. Despite the simple network architecture, the performance of this network is very strong, with an AUC of over 0.99. A zero-background, high-efficiency region can be constructed through cuts of around 0.8 on the output score. While more complex models could likely improve on these results, this already represents near-perfect classification, and will serve as a basis for comparison with other network structures.
Figure 7: Results of training the fully-connected network with a BCE loss function and the Gaussian dataset. The output scores (left) show good separation of signal and background for the Gaussian-distributed features in the generated dataset, even with the relatively simple network architecture. The ROC curve and AUC (right) provide a measure of the performance of the network.
The one-to-one linear network is trained in the same way, using BCE loss, with results shown in Fig. 8. The classifier is not as strong, since this model is equivalent to applying cuts on all features such that signal events failing any one cut are misclassified, while the simple linear network can compensate for background-like behavior in one feature with signal-like behavior in other features. The biases of the one-to-one linear network can be translated into cuts on the input features, shown in Fig. 9. The signal efficiency (true positive rate) of these cuts applied to the test sample is 93%, while the background efficiency (false positive rate) is 5%.
Figure 8: The output scores (left), ROC curve and AUC (right) after training the one-to-one linear network with a BCE loss function and the Gaussian dataset.
Figure 9: Gaussian dataset features for signal (red) and background (blue) shown with cuts optimized using a BCE loss function. Dashed lines represent "greater-than" cuts and dotted lines represent "less than" cuts. The cuts have a signal efficiency of 93% and a background efficiency of 3%.
The one-to-one linear model is next trained with the efficiency loss rather than BCE loss, where we set the target efficiency to be 93% to match the output of the previous configuration. After training, the cuts extracted using the efficiency loss give an actual signal efficiency of 92% with a background efficiency of 2.7%, with nearly identical cuts to the network trained with BCE loss. The choice of hyperparameters $\alpha$ and $\beta$, which control the importance of reaching the target signal efficiency and minimal background efficiency, respectively, influences how close the network gets to the target signal efficiency before the background efficiency becomes too large. In this case, the target signal efficiency is not reached, in favor of keeping the background efficiency small. Some tuning of the $\alpha$ and $\beta$ parameters will likely be needed for other applications depending on the composition of the dataset and the desired signal efficiency or background rejection.
Finally, an ensemble of one-to-one linear networks trained with the contractive efficiency loss function to create cuts that target signal efficiencies from 20% to 90% in steps of 10%. Additional models for target signal efficiencies of 93% and 95% are also included, to investigate the regime where background rejection and signal efficiency begin to compete in the minimization of the loss.
The loss distribution for the efficiency scan network is shown in Fig. 10, including the contributions from the different terms in the training loss. The loss terms associated with reaching the desired signal efficiency and minimizing the background efficiency dominate the overall loss, but the smaller terms associated with the size of the cuts and the smoothness of cuts across target efficiencies regulate the behavior and prevent the cuts from settling into inappropriately loose or tight values. The impact of the BCE term is negligible, and removing it has no impact on the derived cuts in this fit configuration.
Figure 10: The loss distribution for the efficiency scan network, for training and test subsets of the Gaussian dataset. The loss is composed of different terms, corresponding to the signal efficiency, background efficiency, cut sizes, and smoothness conditions, along with a BCE term, all scaled by independent hyperparameters listed in Table 1.
The output score distributions and ROC curves for the trained networks are shown in Figs. 11 and 12. The ROC curves show that the networks trained for low signal efficiencies are far from optimal as measured by AUC. The score distributions show that the signal peaks near zero and near one, with events migrating from zero to one as the target efficiency rises. Due to the way the nets are trained, the vast majority of events passing the trained cuts have output scores above 0.5, and all events that fail cuts have output scores less than 0.5. Taken together, this implies that for a given target efficiency, the effective cut on the output score distribution that corresponds to an application of the trained cuts is near 0.5, in a part of the output score distribution that is not rapidly changing. This makes the calculation of the output score less sensitive to variations in the input features that might lead to changes in the output score. For comparison, the output score corresponding to 40% (50%) for the network with BCE loss is 0.9999994 (0.9999969), so changes of a few parts in $10^{-6}$ is enough to change the efficiency in such a network.
Figure 11: Output score distributions for different target efficiencies in the efficiency scan ensemble trained using the Gaussian dataset.
Figure 12: ROC curves for different target efficiencies in the efficiency scan ensemble trained using the Gaussian dataset.
Figure 13 shows the evolution of cut values across different target efficiencies, with a comparison to the BCE-trained cuts at an efficiency of 93%. The cuts vary smoothly across efficiencies, as desired, and closely match the cuts for the BCE-trained net. The tension between reaching the target signal efficiency and minimizing the background efficiency, especially at large target efficiencies, is shown in Fig. 14.
@@ 6.1 Gaussian dataset
Figure 13: Cuts on Gaussian dataset features extracted from the trained one-to-one linear networks in the efficiency scan ensemble, as a function of the target efficiency. The cuts from the BCE-trained network are shown for comparison.
Figure 14: Signal and background efficiencies as a function of the target efficiency for the Gaussian dataset, both calculated by the efficiency loss function and after applying the cuts post-training, for subnets in ensemble of one-to-one linear networks.
6.2 SUSY dataset
The set of SUSY and SM $WW$ events is used to study the CABIN approach in a more realistic context for collider physics, in which the collection of features does not provide perfect discrimination between signal and background. The seven features described in Section 5 are used as inputs. Results of training the simple fully-connected network on the SUSY dataset using a BCE loss function are shown in Fig. 15. This configuration is the only case in which the learning rate was changed from 0.5 to 0.1, to avoid oscillations observed in the loss when near the minimum. The simple fully-connected network reaches an AUC of 0.69, reflecting the increased difficulty of this classification task relative to classification in the Gaussian dataset. [^1] For comparison, a network with two hidden layers and 20 nodes/layer achieves an AUC of 0.85, on the same dataset. This is further seen in Figs. 16 and 17, where the efficiency scan network is able to achieve higher ROC scores than the fully-connected network for all target efficiencies. In all cases, the score distribution is bimodal, reflecting the binary pass/fail decisions the network is modeling.
Figure 15: Results of training the fully-connected network with the SUSY dataset.
Figure 16: Output score distributions for different target efficiencies in the efficiency scan ensemble when trained with SUSY dataset.
Figure 17: ROC curves for different target efficiencies in the efficiency scan ensemble trained using the SUSY dataset.
@@ 6.2 SUSY dataset
The loss distribution for the efficiency scan network trained on the SUSY dataset is shown in Fig. 18, including the contributions from the different terms in the training loss. The loss terms associated with reaching the desired signal efficiency and minimizing the background efficiency dominate the overall loss, but the smaller terms associated with the size of the cuts and the smoothness of cuts across target efficiencies regulate the behavior and prevent the cuts from settling into inappropriately loose or tight values. The impact of the BCE term is negligible, and removing it has no impact on the derived cuts in this fit configuration.
@@
The optimized cuts on the SUSY dataset for the efficiency scan network and for a one-to-one linear network trained with a BCE loss function are shown in Fig. 19. The network trained with a BCE loss function has a signal efficiency of 67% and a background efficiency of 41%, and fails to converge on cuts that perform as well as the efficiency scan network. This indicates that the smoothing terms in the loss function do not always lead to worse performance, and can help to avoid convergence to a local minimum that the learning routine cannot escape. The efficiency scan network also succeeds at ensuring a smooth (but still nonlinear) evolution of cut values for the individual input features in the SUSY dataset.
Figure 19: Cuts on SUSY dataset features extracted from the trained one-to-one linear networks in the efficiency scan ensemble, as a function of the target efficiency. The cuts from the BCE-trained network is shown for comparison.
The correspondence between the approximated efficiencies computed in the loss function and the actual efficiencies of the trained cuts is shown in Fig. 20. Here it is apparent that the background efficiency rises proportionally with the signal efficiency, rather than remaining zero for all but the highest target signal efficiencies as with the Gaussian dataset. The choice of a final working point for an analysis would depend on factors such as the statistical and systematic uncertainties on the background yields and the signal cross section.
Figure 20: Signal and background efficiencies, calculated in the efficiency loss function and after applying the cuts post-training, as a function of the target efficiency for subnets in ensemble of one-to-one linear networks trained using the SUSY dataset.
@@ 7 Discussion
To fill this gap in optimization tools, we present the CABIN approach to learning optimized cuts for event classification, where the cut values are represented by biases in a type of linear neural network. The CABIN technique uses gradient-based learning and is aware of useful HEP metrics such as signal and background efficiency. The method can self-regulate cuts across working points to deliver both good performance and robustness against statistical noise. An implementation of this approach is available in the cabin python package [cabin].
Extensions of this approach can be used to optimize cuts based on significance metrics [ATL-PHYS-PUB-2020-025], possibly with additional requirements to ensure signal yield large enough to enable discovery or limit setting. Another foreseen extension will be ensuring the smooth evolution cuts as functions of particle or event properties that cause correlated changes in input features. For example, when training a classifier for particle identification, the input features may vary with particle energy, and will likely evolve monotonically, though possibly in different ways for signal and background. Contractive terms in the loss function can ensure that cuts on input features for neighboring energy bins have similar values, consistent with the expected smooth evolution of the features themselves.
Finally, as suggested in Ref. [Watts:2024wes], a CABIN network can be embedded as one step in a chain of differentiable analysis selections. This provides a path towards full-chain optimization of cut-based event generation, triggering, and analysis pre-selection, followed by final optimization using fully connected networks, all within the same minimization task.

### v2

@@ 5 Results
The ROC curves for the methods above applied to the SUSY data sample are shown in Fig. 6. The BDT provides the best performance as measured by the summary AUC scores. The CABIN efficiency scan, TMVA kCuts, and simple fully-connected neural network all achieve equivalent AUC scores, though the ROC curves demonstrate that the methods have different performance depending on the target efficiency. The performance of a set of cuts derived using a one-to-one linear network with a BCE loss function are also shown for comparison.
Figure 6: Summary of ROC curves and AUC scores for different methods applied to the SUSY dataset. For the CABIN and TMVA methods, true-positive and false-negative rates are assessed after applying the cuts derived from those methods. ROC curves for the BDT and Simple Neural Network methods are derived from the distribution of output scores.
The optimized cuts for the SUSY dataset from the efficiency scan network, from TMVA kCuts, and from a one-to-one linear network trained with a BCE loss function are shown in Fig. 7. The efficiency scan network succeeds at ensuring a smooth (but still nonlinear) evolution of cut values for the individual input features in the SUSY dataset, while the cuts tuned by TMVA kCuts vary significantly as a function of the target efficiency.
@@
Figure 7: Cuts on SUSY dataset features extracted from the trained one-to-one linear networks in the efficiency scan ensemble, and from TMVA kCuts, as a function of the target efficiency. The cuts from the BCE-trained network are also shown for comparison. Solid lines indicate "greater-than" cuts, while dashed lines indicate "less-than" cuts. The cut on $\Delta\phi(\ell_{2},E_{\mathrm{T}}^{\mathrm{miss}})$ is the only one where kCuts and CABIN have opposite orientation.
@@
The loss distribution for the efficiency scan network trained on the SUSY dataset is shown in Fig. 8, including the contributions from the different terms in the training loss. The loss terms associated with reaching the desired signal efficiency and minimizing the background efficiency dominate the overall loss, but the smaller terms associated with the size of the cuts and the smoothness of cuts across target efficiencies regulate the behavior and prevent the cuts from settling into inappropriately loose or tight values. The impact of the BCE term appears large, but removing it has no impact on the derived cuts.
@@ 6 Discussion
To fill this gap in optimization tools, we present the CABIN approach to learning optimized cuts for event classification, where the cut values are represented by biases in a type of linear neural network. The CABIN technique uses gradient-based learning and is aware of useful HEP metrics such as signal and background efficiency. The method can self-regulate cuts across working points to deliver both good performance and robustness against statistical noise. An implementation of this approach is available in the cabin python package [21].
Extensions of this approach can be used to optimize cuts based on significance metrics [25], possibly with additional requirements to ensure signal yield large enough to enable discovery or limit setting. Another foreseen extension will be ensuring the smooth evolution cuts as functions of particle or event properties that cause correlated changes in input features. For example, when training a classifier for particle identification, the input features may vary with particle energy, and will likely evolve monotonically, though possibly in different ways for signal and background. Contractive terms in the loss function can ensure that cuts on input features for neighboring energy bins have similar values, consistent with the expected smooth evolution of the features themselves.
Finally, as suggested in Ref. [12], a CABIN network can be embedded as one step in a chain of differentiable analysis selections. This provides a path towards full-chain optimization of cut-based event generation, triggering, and analysis pre-selection, followed by final optimization using fully connected networks, all within the same minimization task.